# Multi-Server FL with Overlapping Clients: A Latency-Aware Relay Framework

### v1

@@ I Introduction
Federated Learning (FL) has gained considerable attention due to its ability to enable collaborative model training while preserving data privacy [1, 2]. A typical FL system operates in rounds, where clients perform local training based on their own datasets and then upload model updates to a cloud server for aggregation and broadcasting in the next round of training. However, the limitation of this single-server architecture is the long-range communication latency between clients and the cloud server (CS), especially in large-scale FL systems. To address this issue, multi-server FL architectures have been proposed [3, 4]. A prominent example is hierarchical federated learning (HFL) [3], where ESs are responsible for aggregating local model updates from their associated clients and subsequently uploading their aggregated models to the CS. However, frequent communications between ESs and the CS can still lead to high latency and resource consumption.
In [5] and [6], a new FL architecture utilizing multiple servers is studied, which exploits the realistic deployment of 5G-and-beyond networks where a client can be located in the overlapping coverage areas of multiple servers. The key idea is that clients download multiple models from all the edge servers they can access and train their local models based on the average of these models. However, OCs must wait for all relevant ESs to deliver their models before training, leading to considerable delay when inter-edge heterogeneity is pronounced. Moreover, when the number of OCs is limited, their contributions to global model convergence is marginal.
To solve these challenges, we have proposed FedOC [7], where OCs can act as relay nodes that forward edge models between adjacent cells upon reception, thereby enabling distributed aggregation across all ESs. However, this work proved convergence only for three cells, and each ES model in the current round could reach only its neighboring ESs. In this paper, we extend the convergence proof to an arbitrary number of cells and optimize ES transmission timing to maximize the propagation range of each ES model among all ESs, thereby accelerating convergence within a finite time horizon.
@@
  • We propose a distributed aggregated multi-server FL framework, where we exploit the OCs for model exchange between neighboring cells and facilitate cross-server collaboration. We derive the upper bound of the gap between the expected and optimal global loss values with respect to the propagation range of each ES model among all ESs. Based on this, we merge model relaying with local update uploading into a single transmission, and further optimize the model relay strategy and transmission scheduling to neighboring ESs for each ES model, so as to maximize the propagation range of each ES model within a finite time without introducing new communication links.
  • To solve the optimization problem, we design a conflict-graph-based local search algorithm that jointly considers path dependency and communication constraints. For each ES $l$, the algorithm identifies both the farthest reachable path and the farthest received model path, constructing a conflict graph to capture mutual transmission conflicts. By combining greedy initialization with iterative local search, it maximizes the path weight, i.e., the total aggregated data volume across all ESs, thereby enhancing overall model propagation depth.
  • We conduct extensive experiments on multiple datasets under varying numbers of cells and heterogeneous data distribution. The results demonstrate that our algorithm consistently outperforms the compared benchmarks in terms of both convergence accuracy and speed.
II System Model
II-A Network Architecture
@@
As shown in Fig. 1, we model overlaps using a chain topology, where each ES $l$ overlaps with its immediate neighbors, ES l1l-1 and $l+1$, except for boundary ESs (ES 1 and ES $L$), which overlap with only one neighbor. While not necessarily implying a physical line, this structure captures typical cases of localized overlaps, e.g., due to limited signal range or linear deployment patterns like roads or urban blocks. It simplifies analysis while preserving the key challenges posed by overlapping coverage and multi-server coordination [7]. A single client in each overlap region, denoted $b_{l,l+1}$, serves as the Relay Overlapping Client (ROC) to forward models between ESs, while the rest are Normal OCs (NOCs). The commonly used notations are summarized in Table I.
@@ III THEORETICAL RESULTS
IV Problem Formulation and Solution
@@
V Performance Evaluation
V-A Simulation Setup
@@ VI Conclusion
Under these definitions, each ROC is mathematically associated with a unique cell. As shown in (II-B) the edge model of each ES $l$ is the weighted sum of all participating ES models.
@@
We next focus on bounding $A_{1}$. Based on the definitions of $\bm{w}_{R}^{(f_{l})}\text{ and }\bm{w}_{R}^{(c)}$ in (II-B) and (19), we have
@@
where equality 3 holds because (17) and (18). Inequality 4 holds because the Assumption 1 and we denote $g_{\max}\left(\bm{w}_{r,q}^{(c_{j})}\right)=\max_{i=1}^{C}\left\|\nabla_{\bm{w}}\mathbb{E}_{\bm{x}\mid y=i}\log f_{i}\left(\bm{x},\bm{w}_{r,q}^{(c_{j})}\right)\right\|$ for $r\in\{1,2,\ldots,R\}\text{ and }q\in\{0,1,\ldots,E\}$. equality 5 holds because we denote $a_{r,q}^{(k)}=1+\eta_{r,q}\sum_{i=1}^{C}P_{y=i}^{(k)}\lambda_{\bm{x}\mid y=i}$ for $r\in\{1,2,\ldots,R\}\text{ and }q\in\{0,1,\ldots,E\}$. From (VI) we deduce the relation between $\left\|\bm{w}_{R-1,E-1}^{(k)}-\bm{w}_{R-1,E-1}^{(c_{j})}\right\|\text{ and }\left\|\bm{w}_{R-1,E-2}^{(k)}-\bm{w}_{R-1,E-2}^{(c_{j})}\right\|$.
@@
where we assume $\prod_{a}^{b}a_{r,e}^{(k)}=1,~\forall b<a,~r~\text{and}~e$. Institute (VI) into (VI), we have
@@
where $D_{R-1}^{(j)}=\frac{\hat{N}_{R-1}^{(f_{j})}}{\sum_{j=1}^{L}\hat{N}_{R-1}^{(f_{j})}}\sum_{k=1}^{\hat{K}_{r}^{(f_{j})}}\frac{n^{(k)}}{\hat{N}_{R-1}^{(f_{j})}}\left(\prod_{e=0}^{E-1}a_{R-1,e}^{(k)}\right)$, and $G_{R-1}^{(j)}=\frac{\hat{N}_{R-1}^{(f_{j})}}{\sum_{j=1}^{L}\hat{N}_{R-1}^{(f_{j})}}\sum_{k=1}^{\hat{K}_{r}^{(f_{j})}}\frac{n^{(k)}}{\hat{N}_{R-1}^{(f_{j})}}\left(\sum_{i=1}^{C}\bigl|P_{y=i}^{(k)}-P_{y=i}^{(c_{j})}\bigr|\right)\\ \sum_{e=0}^{E-2}\eta_{R-1,d}\left(\prod_{d=e+1}^{E-1}a_{R-1,d}^{(k)}\right)g_{\max}\!\left(\bm{w}_{R-1,e}^{(c_{j})}\right)$. Plug (VI) into (VI), we have
@@
We assume that there exist constants $g_{\max}$ and $\beta_{r}^{(j)}$ such that $g_{\max}\big(\bm{w}_{r,e}^{(c_{j})}\big)\leq g_{\max}$ and $\beta_{r,e}^{(j)}\leq\beta_{r}^{(j)}$, $\forall r,~e,~j$. Moreover, we denote $\sum_{j=1}^{L}\beta_{r}^{(j)}=\beta_{r}$. Since $a_{r,q}^{(k)}=1+\eta_{r,q}\sum_{i=1}^{C}P_{y=i}^{(k)}\lambda_{\bm{x}\mid y=i}>1,~\forall r,~q$, based on (VI), we further assume $\beta_{\text{min}}\leq\beta_{r}\leq\beta_{\text{max}}$. Then we have
@@
Substituting (VI), (30) and (VI) into (VI), we have

### v2

@@ II Introduction
Federated Learning (FL) has gained considerable attention due to its ability to enable collaborative model training while preserving data privacy [1]. A typical FL system operates in rounds, where clients perform local training based on their own datasets and then upload model updates to a cloud server for aggregation and broadcasting in the next round of training. However, the limitation of this single-server architecture is the long-range communication latency between clients and the cloud server (CS), especially in large-scale FL systems. To address this issue, multi-server FL architectures have been proposed [2, 3]. A prominent example is hierarchical federated learning (HFL) [2], where ESs are responsible for aggregating local model updates from their associated clients and subsequently uploading their aggregated models to the CS. However, frequent communications between ESs and the CS can still lead to high latency and resource consumption.
In [4] and [5], a new FL architecture utilizing multiple servers is studied, which exploits the realistic deployment of 5G-and-beyond networks where a client can be located in the overlapping coverage areas of multiple servers, a situation that is particularly common in dense or mesh-like network deployments. The key idea is that clients download multiple models from all the edge servers they can access and train their local models based on the average of these models. Aggregating multiple edge models mainly enhances each OC’s own generalization. Once uploaded to the ES, its contribution to cell-level generalization is limited by the small data size of each client, thus, the global model performance decreases further as the number of OCs drops. To address these challenges, a few studies have begun exploring OC-assisted cross-cell model sharing [6]. In [6], OCs buffer models from multiple ESs and upload them together after completing local training in the next round, which increases shared information but causes stale updates and higher storage overhead. Building on these insights, we have proposed FedOC [7], where OCs serve as relay nodes that forward edge models between adjacent cells upon reception. Compared with [4] and [5], every client can train on aggregated models from multiple edges in each round, which further enhances generalization and makes the model performance independent of the number of OCs. However, the theoretical results in [7] indicate that, as the number of cells increases, relying solely on a limited number of neighboring ES models makes it difficult for each ES to achieve fast synchronization across the network. In this paper, we propose a new approach. Instead of immediately taking actions, witnin a limited time budget, each ES introduces an appropriate waiting period to aggregate its own model with models received from neighboring ESs before forwarding the aggregated result via OCs. Meanwhile, since OCs, as clients, are required to upload their locally trained models to ESs, we aggregate the relayed ES model with the locally uploaded model and forward the aggregated model to neighboring ESs in OCs. In this way, wider model dissemination among ESs can be achieved without introducing any additional communication links or overhead. Motivated by this idea, we derive an upper bound on the optimality gap of the loss function between the proposed distributed aggregation scheme and the optimal solution. Our theoretical analysis shows that the global model becomes closer to the optimal solution as each ES receives more models in the current round. However, waiting for more ES models to be relayed through particular OCs introduces additional latency at each ES, thereby leading to a fundamental tradeoff between model performance and communication latency. Within the framework of leveraging OCs for inter-ES model relaying, we dynamically optimize the model transmission timing of each ES to maximize the number of ES models received by each ES within a finite time period. Without introducing any additional communication links or requiring synchronization with a cloud center, this design promotes distributed model fusion.
@@
  • We propose a distributed aggregated multi-server FL framework, where we exploit OCs as relays for model exchange between neighboring cells and facilitate cross-server collaboration. To reduce relay communication overhead, we assign a single relay OC per overlapping region and merge model relaying with local update uploading into a single transmission. We derive the upper bound of the gap between the expected and optimal global losses with respect to the propagation range of each ES model among all ESs. Based on this, we optimize the model relay strategy and transmission scheduling to neighboring ESs for each ES model, so as to maximize the propagation range of each ES model within a finite time without introducing new communication links.
  • To solve the optimization problem, we design a conflict-graph-based local search algorithm that jointly optimizes transmission decisions and durations among ESs. By constructing a conflict graph to model transmission interference and applying a greedy–local search strategy, the algorithm maximizes the transmission depth of each ES model among all ESs in a cloud-free manner, thereby improving model propagation efficiency.
  • We conduct extensive experiments on multiple datasets under varying numbers of cells and heterogeneous data distribution. The results demonstrate that our algorithm consistently outperforms the compared benchmarks in terms of both convergence and accuracy.
III System Model
III-A Network Architecture
@@
As shown in Fig. 1, we model overlaps using a chain topology, where each ES $l$ overlaps with its immediate neighbors, ES l1l-1 and $l+1$, except for boundary ESs (ES 1 and ES $L$), which overlap with only one neighbor. The chain topology is adopted for two main reasons. First, it simplifies analysis while preserving the key challenges posed by overlapping coverage and multi-server coordination. Second, it can be readily extended to more complex multi-server topologies, e.g., dense mesh deployments where one ES overlaps with three or more neighboring ESs. When a single region is simultaneously overlapped by more than three cells, severe communication interference arises, and the available bandwidth for each cell becomes further fragmented, resulting in substantial communication delays. By limiting each overlapped region to at most two logically participating ESs per communication round, a chain-like logical topology transforms complex mesh structures into simple linear subgraphs, enabling efficient operation without requiring additional complex interference management or communication coordination mechanisms. A single client in each overlap region, denoted $b_{l,l+1}$, serves as the Relay Overlapping Client (ROC) to forward models between ESs, while the rest are Normal OCs (NOCs). The commonly used notations are summarized in Table I.
@@ IV THEORETICAL RESULTS
V Problem Formulation and Solution
@@
VI Performance Evaluation
VI-A Simulation Setup
@@ VII Conclusion
Notice that $\bm{w}_{r,E}^{(l-1,l)}$ in (3) may contain not only the locally aggregated model of cell l1l-1 but also models relayed from other cells through multi-hop transmissions. For instance, the model of cell $i$ can be forwarded along the path $\mathcal{P}_{i\rightarrow(l-1)}=\{\text{ES }i,\text{ROC }b_{i,i+1},\text{ES }(i+1),\dots,\text{ROC }b_{l-2,l-1},\text{ES }(l-1)\}$, where each relay node aggregates the received model with its own model before forwarding, provided that the timing constraint is satisfied. We denote $\bm{p}_{r}^{(f_{l})}=[p_{r}^{(1,l)},\ldots,p_{r}^{(L,l)}]$, where $p_{r}^{(i,l)}\in\{0,1\}$ indicates whether cell $i$’s model participates in ES $l$’s aggregation. Based on (3), we have:
$\quad+p_{r}^{(l,l)}\tilde{N}_{r}^{(f_{l})}\bm{\tilde{w}}_{r,E}^{(f_{l})}$
where $\bm{\hat{w}}_{r,E}^{(f_{i})}$ is denoted as
@@
Under these definitions, each ROC is mathematically associated with a unique cell. As shown in (VII) the edge model of each ES $l$ is the weighted sum of all participating ES models.
@@
We next focus on bounding $A_{1}$. Based on the definitions of $\bm{w}_{R}^{(f_{l})}\text{ and }\bm{w}_{R}^{(c)}$ in (VII) and (22), we have
@@
where equality 3 holds because (20) and (21). Inequality 4 holds because the Assumption 1 and we denote $g_{\max}\left(\bm{w}_{r,q}^{(c_{j})}\right)=\max_{i=1}^{C}\left\|\nabla_{\bm{w}}\mathbb{E}_{\bm{x}\mid y=i}\log f_{i}\left(\bm{x},\bm{w}_{r,q}^{(c_{j})}\right)\right\|$ for $r\in\{1,2,\ldots,R\}\text{ and }q\in\{0,1,\ldots,E\}$. equality 5 holds because we denote $a_{r,q}^{(k)}=1+\eta_{r,q}\sum_{i=1}^{C}P_{y=i}^{(k)}\lambda_{\bm{x}\mid y=i}$ for $r\in\{1,2,\ldots,R\}\text{ and }q\in\{0,1,\ldots,E\}$. From (VII) we deduce the relation between $\left\|\bm{w}_{R-1,E-1}^{(k)}-\bm{w}_{R-1,E-1}^{(c_{j})}\right\|\text{ and }\left\|\bm{w}_{R-1,E-2}^{(k)}-\bm{w}_{R-1,E-2}^{(c_{j})}\right\|$.
@@
where we assume $\prod_{a}^{b}a_{r,e}^{(k)}=1,~\forall b<a,~r~\text{and}~e$. Institute (VII) into (VII), we have
@@
where $D_{R-1}^{(j)}=\frac{\hat{N}_{R-1}^{(f_{j})}}{\sum_{j=1}^{L}\hat{N}_{R-1}^{(f_{j})}}\sum_{k=1}^{\hat{K}_{r}^{(f_{j})}}\frac{n^{(k)}}{\hat{N}_{R-1}^{(f_{j})}}\left(\prod_{e=0}^{E-1}a_{R-1,e}^{(k)}\right)$, and $G_{R-1}^{(j)}=\frac{\hat{N}_{R-1}^{(f_{j})}}{\sum_{j=1}^{L}\hat{N}_{R-1}^{(f_{j})}}\sum_{k=1}^{\hat{K}_{r}^{(f_{j})}}\frac{n^{(k)}}{\hat{N}_{R-1}^{(f_{j})}}\left(\sum_{i=1}^{C}\bigl|P_{y=i}^{(k)}-P_{y=i}^{(c_{j})}\bigr|\right)\\ \sum_{e=0}^{E-2}\eta_{R-1,d}\left(\prod_{d=e+1}^{E-1}a_{R-1,d}^{(k)}\right)g_{\max}\!\left(\bm{w}_{R-1,e}^{(c_{j})}\right)$. Plug (VII) into (VII), we have
@@
We assume that there exist constants $g_{\max}$ and $\beta_{r}^{(j)}$ such that $g_{\max}\big(\bm{w}_{r,e}^{(c_{j})}\big)\leq g_{\max}$ and $\beta_{r,e}^{(j)}\leq\beta_{r}^{(j)}$, $\forall r,~e,~j$. Moreover, we denote $\sum_{j=1}^{L}\beta_{r}^{(j)}=\beta_{r}$. Since $a_{r,q}^{(k)}=1+\eta_{r,q}\sum_{i=1}^{C}P_{y=i}^{(k)}\lambda_{\bm{x}\mid y=i}>1,~\forall r,~q$, based on (VII), we further assume $\beta_{\text{min}}\leq\beta_{r}\leq\beta_{\text{max}}$. Then we have
@@
Substituting (VII), (33) and (VII) into (VII), we have